OpenStudy (anonymous):

Find the equation of the line tangent to y = –x2 + 3x + 8 at x = 2.

OpenStudy (anonymous):

@dan815 could you help me?

OpenStudy (anonymous):

the derivative formula that I got is f'(x)= -6x+1

OpenStudy (anonymous):

but none of the answers match my answer

OpenStudy (anonymous):

Find the equation of the line tangent to y = –x2 + 3x + 8 at x = 2. y = –2x + 2 y = –x + 12 y = –x + 2 y = –2x – 12

OpenStudy (dan815):